# Continuum model for the terahertz dielectric response of glasses

### v1

@@ Abstract
where $\tilde{G}(\omega)=G(\omega)/\rho$ and $\tilde{M}(\omega)=M(\omega)/\rho$. Equation (4) is the expression used to analyze the THz dielectric response and the BP-related excess contribution. For quantitative comparison with experiment, the model inputs are $\rho$, $k_{D}$, $G(\omega)$, $K$, and $\Delta q(k)$. For glycerol we use an effective shear modulus $G(\omega)$ obtained from a heterogeneous-elasticity-theory coherent-potential-approximation (HET-CPA) analysis of the vibrational density of states $g(\omega)$, following our previous work [44], and parameterize the IR-effective charge spectrum by retaining the leading terms of a Maclaurin expansion, $\Delta q(k)\approx q_{0}+q_{2}k^{2}$. Because the dielectric spectra are insensitive to $q_{0}$ within the experimental uncertainty, we fix $q_{0}=0$. The remaining parameter $q_{2}$ is selected by a parameter scan to minimize the misfit between calculated and measured $\varepsilon^{\prime}(\omega)$ and $\varepsilon^{\prime\prime}(\omega)$ (see Supplemental Material [38]). The parameters used in Fig. 1 are summarized in Table 1.
@@
In Eq. (4) all microscopic information on the IR coupling is collected into the $k$-dependent oscillator strength $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$. Since the $k$ integral is restricted to the acoustic window $0<k<k_{\rm D}$, we only require $\Delta q(k)$ over this $k$ range. Because the quantity entering Eq. (4) is the isotropic disorder-averaged oscillator strength $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$, which is expected to be a smooth (analytic) scalar function at $k\to 0$, we adopt the minimal quadratic form $\Delta q(k)\approx q_{0}+q_{2}k^{2}$ and treat $q_{0}$ and $q_{2}$ as phenomenological parameters. Microscopically, however, the IR oscillator strength of each normal mode is determined by the mode-carried charge fluctuations, equivalently by the charge–charge dynamic structure factor $S_{ZZ}(k,\omega)$ and its projection onto the eigenmodes [48, Chap. 10]. A fully microscopic determination of $\Delta q(k)$ would thus require an explicit evaluation of $S_{ZZ}(k,\omega)$, which is beyond the scope of the present work. Here we use the static structure factor $S_{ZZ}(k)$ from molecular dynamics (MD) [49] only as a qualitative diagnostic of medium-range charge correlations on the length scale associated with the first sharp diffraction peak (FSDP). For glycerol, $S_{ZZ}(k)$ exhibits no maximum at the FSDP ($k\sim 1.5~\text{\AA }^{-1}$ [50]), and its lowest-$k$ maximum occurs at the interatomic-spacing scale (Fig. 3). Because the relevant $k$ window ($0<k<k_{\rm D}$) lies on the low-$k$ side of this maximum, the static $S_{ZZ}(k)$ varies smoothly over the range contributing to Eq. (4). Accordingly, we do not introduce any explicit FSDP-like structure into the phenomenological $\Delta q(k)$.
@@
Finally, we turn to the question of why Taraskin’s model does not reproduce the nearly linear $C_{\rm IR}(\omega)$ observed near the BP. Taraskin’s approach starts from a factorization ansatz $\alpha(\omega)\approx C_{\rm IR}(\omega)g(\omega)$ [35], motivated by the mode-resolved linear-response expression for crystals derived by Maradudin [34]. Microscopically, however, linear response gives $\alpha(\omega)=\sum_{\kappa}C_{\kappa}\delta(\omega-\omega_{\kappa})$, i.e., a sum over modes with mode-dependent couplings $C_{\kappa}$, not a product of a smooth coupling and the VDOS $g(\omega)=\sum_{\kappa}\delta(\omega-\omega_{\kappa})$. Identifying $\alpha(\omega)$ with $C_{\rm IR}(\omega)g(\omega)$ amounts to a modeling assumption: it replaces the mode-resolved couplings $C_{\kappa}$ by a single smooth function of frequency $C_{\rm IR}(\omega)$. This limitation was emphasized early on for disordered solids, where absorption involves mode- (or subband-) resolved couplings that can vary strongly with frequency (see Galeener and Sen [51]). A related non-factorizable convolution form appears in low-frequency Raman scattering (Schmid and Schirmacher [52]). Under Taraskin’s decomposition into uncorrelated and correlated charge components, this factorization yields the quadratic form $C_{\rm IR}(\omega)=A+B\omega^{2}$ used in their analysis. Within our framework, Eq. (4) can be interpreted as a continuum $k$-space coarse graining of Maradudin’s mode-resolved linear-response formula [34], in which the discrete mode sum is replaced by an integral over $k$ and the mode-dependent couplings are represented phenomenologically by an effective $k$-dependent oscillator strength $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$. In this coarse-grained picture, the spectral factor is the viscoelastic Green’s function of the acoustic branches, $\left[-\omega^{2}+\tilde{C}_{\alpha}(\omega)k^{2}\right]^{-1}$, determined by the complex moduli $G(\omega)$ and $M(\omega)$. A fully microscopic determination of these effective inputs is left for future work.
@@ Acknowledgements.
For the charge–charge correlations, we use the same partial functions but weight them by the (dimensionless) partial charges $z_{\alpha}$ (i.e., the atomic partial charge is $ez_{\alpha}$). We follow the conventional multicomponent definition of the charge–charge structure factor [48, Chap. 10] and use the normalized form:
@@
S10. Rescaling procedure for the $R$-scan in Fig. 4
@@
Figure S2: (a) Real and imaginary parts of the effective shear modulus $G(\omega)$ from the HET-CPA analysis, plotted as $\mathrm{Re}[G(\omega)/G_{0}]$ (orange) and $\mathrm{Im}[G(\omega)/G_{0}]$ (blue), with $G_{0}=5.43~\mathrm{GPa}$. (b) Reduced vibrational density of states $g(\omega)/(A_{\mathrm{D}}\omega^{2})$ computed from $G(\omega)$ (magenta line) compared with INS data [S2] (filled circles).

### v2

@@ Abstract
where $\tilde{G}(\omega)=G(\omega)/\rho$ and $\tilde{M}(\omega)=M(\omega)/\rho$. Equation (4) is the expression used to analyze the THz dielectric response and the BP-related excess contribution. For quantitative comparison with experiment, the model inputs are $\rho$, $k_{\rm D}$, $G(\omega)$, $K$, and $\Delta q(k)$. For glycerol we use an effective shear modulus $G(\omega)$ obtained from a heterogeneous-elasticity-theory coherent-potential-approximation (HET-CPA) analysis of the vibrational density of states $g(\omega)$, following our previous work [47], and parameterize the IR-effective charge spectrum by retaining the leading terms of a Maclaurin expansion, $\Delta q(k)\approx q_{0}+q_{2}k^{2}$. Because the dielectric spectra are insensitive to $q_{0}$ within the experimental uncertainty, we fix $q_{0}=0$. The remaining parameter $q_{2}$ is selected by a parameter scan to minimize the misfit between calculated and measured $\varepsilon^{\prime}(\omega)$ and $\varepsilon^{\prime\prime}(\omega)$ (see Supplemental Material [38]). The parameters used in Fig. 1 are summarized in Table 1.
@@
In Eq. (4) all microscopic information on the THz (IR) charge coupling is collected into the $k$-dependent oscillator strength $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$. Since the $k$ integral is restricted to the acoustic window $0<k<k_{\rm D}$, we only require $\Delta q(k)$ over this range and adopt the minimal analytic form $\Delta q(k)\approx q_{0}+q_{2}k^{2}$. Microscopically, the mode oscillator strength is set by the mode effective charge $\mathbf{F}^{(p)}$ [51], i.e., the projection of the rigid-ion charge distribution onto each eigenmode $p$ (equivalently by the charge–charge dynamic structure factor $S_{ZZ}(k,\omega)$ and its projection onto the eigenmodes [52, Chap. 10]). In the BP regime, a plane-wave representation of acousticlike eigenmodes with mode-dependent weights $a_{\mathbf{k}\alpha}(\omega_{p})$ [53, 35] makes this projection proportional to the static charge–charge structure factor, $\left\langle\,|\hat{\mathbf{e}}\cdot\mathbf{F}^{(p)}|^{2}\,\right\rangle\propto\sum_{\mathbf{k},\alpha}|a_{\mathbf{k}\alpha}(\omega_{p})|^{2}S_{ZZ}(k)\,\gamma_{\alpha}$ (see Supplemental Material [38]), providing a microscopic route by which a static $S_{ZZ}(k)$ can contribute to the effective coupling spectrum represented by $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$ in Eq. (4). A fully microscopic determination of $\Delta q(k)$ would still require $S_{ZZ}(k,\omega)$, but the static $S_{ZZ}(k)$ from molecular-dynamics (MD) suffices to diagnose whether medium-range correlations (e.g., the first sharp diffraction peak (FSDP)) should appear within the relevant acoustic window. For glycerol, $S_{ZZ}(k)$ exhibits no maximum at the FSDP ($k\simeq 1.5\,{\rm\AA }^{-1}$ [54]) and varies smoothly for $0<k<k_{\rm D}$ (Fig. 3); thus, we do not introduce any explicit FSDP-like structure into $\Delta q(k)$.
@@
Finally, we turn to the question of why Taraskin’s model does not reproduce the nearly linear $C_{\rm IR}(\omega)$ observed near the BP. Taraskin’s approach starts from a factorization ansatz $\alpha(\omega)\approx C_{\rm IR}(\omega)g(\omega)$ [35], motivated by the mode-resolved linear-response expression for crystals derived by Maradudin [34]. Microscopically, however, linear response gives $\alpha(\omega)=\sum_{p}C_{p}\delta(\omega-\omega_{p})$, i.e., a sum over modes with mode-dependent couplings $C_{p}$, not a product of a smooth coupling and the VDOS $g(\omega)=\sum_{p}\delta(\omega-\omega_{p})$. Identifying $\alpha(\omega)$ with $C_{\rm IR}(\omega)g(\omega)$ amounts to a modeling assumption: it replaces the mode-resolved couplings $C_{p}$ by a single smooth function of frequency $C_{\rm IR}(\omega)$. This limitation was emphasized early on for disordered solids, where absorption involves mode- (or subband-) resolved couplings that can vary strongly with frequency (see Galeener and Sen [55]). A related non-factorizable convolution form appears in low-frequency Raman scattering (Schmid and Schirmacher [56]). Under Taraskin’s decomposition into uncorrelated and correlated charge components, this factorization yields the quadratic form $C_{\rm IR}(\omega)=A+B\omega^{2}$ used in their analysis. Within our framework, Eq. (4) can be interpreted as a continuum $k$-space coarse graining of Maradudin’s mode-resolved linear-response formula [34], in which the discrete mode sum is replaced by an integral over $k$ and the mode oscillator strengths are encoded phenomenologically in an effective $k$-dependent spectrum $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$ (see Supplemental Material [38]). In this coarse-grained picture, the spectral factor is the viscoelastic Green’s function of the acoustic branches, $\left[-\omega^{2}+\tilde{C}_{\alpha}(\omega)k^{2}\right]^{-1}$, determined by the complex moduli $G(\omega)$ and $M(\omega)$. A fully microscopic determination of these effective inputs is left for future work.
@@ Acknowledgements.
Section S10 provides a plane-wave expansion and a microscopic route to the static charge–charge structure factor $S_{ZZ}(k)$.
@@
For the charge–charge correlations, we use the same partial functions but weight them by the (dimensionless) partial charges $z_{\alpha}$ (i.e., the atomic partial charge is $ez_{\alpha}$). We follow the conventional multicomponent definition of the charge–charge structure factor [52, Chap. 10] and use the normalized form:
@@
S10. Plane-wave expansion and a microscopic route to the static $S_{ZZ}(k)$
This section provides a microscopic interpretation of the $k$-dependent oscillator-strength spectrum used in the continuum expression for $\varepsilon(\omega)$ in the main text. We start from the standard linear-response mode-sum representation for the vibrational dielectric loss in the harmonic approximation, expressed in terms of mode effective charges. We then show that, when the relevant disordered eigenmodes retain substantial plane-wave character so that a longitudinal/transverse plane-wave expansion provides a useful representation, the disorder- and orientation-averaged mode oscillator strength admits an approximate $k$-space form involving the static charge–charge structure factor $S_{ZZ}(k)$ weighted by the plane-wave coefficients. This provides a microscopic route by which static charge correlations can contribute to the effective coupling spectrum represented phenomenologically by $\langle|\Delta q(k)|^{2}\rangle_{\rm dis}$ in Eq. (4).
(i) Dielectric loss expressed in vibrational eigenmodes.
We recall a standard mode-sum expression for the dielectric loss. For an isotropic system, it can be expressed as[51, 45, 46]:
$\varepsilon^{\prime\prime}(\omega)=\frac{4\pi^{2}}{3V}\sum_{p}\sum_{j=x,y,z}\frac{|F^{(p)}_{j}|^{2}}{2\omega_{p}}\,\delta(\omega-\omega_{p}),$ (S42)
where $p$ labels eigenmodes of frequency $\omega_{p}$ and $F^{(p)}_{j}$ is the Cartesian component of the mode charge-coupling vector (whose squared magnitude gives the oscillator strength). In the most general form (Born effective charges) one has [45, 46]
$F^{(p)}_{j}\equiv\sum_{I,k}\frac{Z_{I,jk}}{\sqrt{m_{I}}}\,e^{(p)}_{I,k}$, (S43)
with $Z_{I,jk}$ the Born effective charge tensor and $\mathbf{e}^{(p)}_{I}$ the (mass-weighted) eigenvector. For the present analysis, we adopt the rigid-ion approximation $Z_{I,jk}\simeq q_{I}\delta_{jk}$, which reduces Eq. (S43) to the mode charge-coupling vector
$\mathbf{F}^{(p)}\equiv\sum_{I}\frac{q_{I}}{\sqrt{m_{I}}}\,\mathbf{e}^{(p)}_{I},\qquad\sum_{j}|F^{(p)}_{j}|^{2}=|\mathbf{F}^{(p)}|^{2}.$ (S44)
Equation (S42) then becomes
$\varepsilon^{\prime\prime}(\omega)=\frac{4\pi^{2}}{3V}\sum_{p}\frac{|\mathbf{F}^{(p)}|^{2}}{2\omega_{p}}\,\delta(\omega-\omega_{p}).$ (S45)
(ii) Plane-wave expansion with explicit longitudinal/transverse polarizations.
In the frequency range where disordered eigenmodes retain substantial plane-wave character, one may expand them in plane waves. Taraskin et al. explicitly state that the low-frequency disordered modes can be approximately expanded in plane waves and that the distribution of the plane-wave coefficients shows two relatively narrow peaks associated with transverse/longitudinal acoustic (TA/LA) hybridization (while broadening at higher frequencies as disorder-induced attenuation becomes stronger) [53, 35]. To make the polarization content explicit, we write
$\mathbf{e}^{(p)}_{I}\simeq\sum_{\mathbf{k}}\sum_{\alpha\in\{L,T_{1},T_{2}\}}\sqrt{\frac{m_{I}}{\bar{m}}}\,a_{\mathbf{k}\alpha}(\omega_{p})\,\hat{\mathbf{e}}_{\alpha}(\mathbf{k})\,e^{i\mathbf{k}\cdot\mathbf{r}_{I}},$ (S46)
where $\bar{m}\equiv N^{-1}\sum_{I}m_{I}$. For each $\mathbf{k}$, the three polarization vectors form an orthonormal triad: $\hat{\mathbf{e}}_{L}(\mathbf{k})\parallel\mathbf{k}$ and $\hat{\mathbf{e}}_{T_{a}}(\mathbf{k})\cdot\mathbf{k}=0$ ($a=1,2$).
(iii) From plane-wave expansion to a static $S_{ZZ}(k)$.
Substituting Eq. (S46) into Eq. (S44) gives
$\mathbf{F}^{(p)}\simeq\frac{1}{\sqrt{\bar{m}}}\sum_{\mathbf{k},\alpha}a_{\mathbf{k}\alpha}(\omega_{p})\,\hat{\mathbf{e}}_{\alpha}(\mathbf{k})\,\rho_{Z}(\mathbf{k}),\qquad\rho_{Z}(\mathbf{k})\equiv\sum_{I}q_{I}e^{i\mathbf{k}\cdot\mathbf{r}_{I}}.$ (S47)
Consider the squared projection onto the unit vector $\hat{\mathbf{e}}$ specifying the direction of the applied electric field. Using statistical isotropy, configurational averaging, and an orientational average over the relative angle between $\hat{\mathbf{e}}$ and the polarization basis $\{\hat{\mathbf{e}}_{\alpha}(\mathbf{k})\}$, cross terms are neglected, yielding the approximate relation
$\left\langle|\hat{\mathbf{e}}\cdot\mathbf{F}^{(p)}|^{2}\right\rangle$ $\simeq\frac{1}{\bar{m}}\sum_{\mathbf{k},\alpha}\left|a_{\mathbf{k}\alpha}(\omega_{p})\right|^{2}\left\langle|\rho_{Z}(\mathbf{k})|^{2}\right\rangle\left\langle(\hat{\mathbf{e}}\cdot\hat{\mathbf{e}}_{\alpha})^{2}\right\rangle_{\rm angles}$
$=\frac{N\bar{q}^{\,2}}{\bar{m}}\sum_{\mathbf{k},\alpha}\left|a_{\mathbf{k}\alpha}(\omega_{p})\right|^{2}S_{ZZ}(k)\,\gamma_{\alpha},$ (S48)
where $\bar{q}^{\,2}\equiv N^{-1}\sum_{I}q_{I}^{2}$, and
$S_{ZZ}(k)\equiv\frac{1}{N\bar{q}^{\,2}}\left\langle\rho_{Z}(\mathbf{k})\rho_{Z}(-\mathbf{k})\right\rangle$
is the normalized static charge–charge structure factor, consistent with the definition adopted in Sec. S8 (note that $q_{I}=ez_{I}$ implies $\bar{q}^{\,2}=e^{2}\bar{z}^{\,2}$ in the notation used there), and
$\gamma_{\alpha}\equiv\left\langle(\hat{\mathbf{e}}\cdot\hat{\mathbf{e}}_{\alpha})^{2}\right\rangle_{\rm angles}.$ (S49)
In an isotropic three-dimensional medium the three polarizations are statistically equivalent, giving $\gamma_{L}=\gamma_{T_{1}}=\gamma_{T_{2}}=1/3$ (hence a transverse-to-longitudinal weight ratio $2:1$). If one further assumes that the two transverse polarizations are equivalent on average, $|a_{\mathbf{k}T_{1}}|^{2}\simeq|a_{\mathbf{k}T_{2}}|^{2}\equiv|a_{\mathbf{k}T}|^{2}$, then Eq. (S48) can be grouped as
$\left\langle|\hat{\mathbf{e}}\cdot\mathbf{F}^{(p)}|^{2}\right\rangle\simeq\frac{N\bar{q}^{\,2}}{\bar{m}}\sum_{\mathbf{k}}\left[\frac{2}{3}\left|a_{\mathbf{k}T}(\omega_{p})\right|^{2}+\frac{1}{3}\left|a_{\mathbf{k}L}(\omega_{p})\right|^{2}\right]\,S_{ZZ}(k).$ (S50)
(iv) Back to $\varepsilon^{\prime\prime}(\omega)$ and comparison with Eq. (4) in the main text.
Using $\frac{1}{3}\sum_{j}|F^{(p)}_{j}|^{2}=\langle|\hat{\mathbf{e}}\cdot\mathbf{F}^{(p)}|^{2}\rangle_{\rm angles}$, Eq. (S45) can be viewed as a mode sum over the orientationally averaged projection. Substituting Eq. (S50) yields the approximate form
$\varepsilon^{\prime\prime}(\omega)\propto\sum_{p}\frac{\delta(\omega-\omega_{p})}{\omega_{p}}\sum_{\mathbf{k}}\left[\frac{2}{3}\left|a_{\mathbf{k}T}(\omega_{p})\right|^{2}+\frac{1}{3}\left|a_{\mathbf{k}L}(\omega_{p})\right|^{2}\right]\,S_{ZZ}(k),$ (S51)
up to an overall prefactor. Equation (S51) makes explicit that (i) a static $S_{ZZ}(k)$ can contribute to the effective oscillator-strength spectrum via the plane-wave expansion, and (ii) the transverse and longitudinal contributions appear in the $2/3$ vs $1/3$ ratio characteristic of an isotropic medium, matching the polarization weighting used in the continuum expression of the main text [Eq. (4)] and in Secs. S3–S5 of this Supplemental Material.
Finally, the physical meaning of the plane-wave weights is analogous to a $k$-space spectral function: Taraskin et al. report that at low frequencies the distribution of $a_{\mathbf{k}\alpha}(\omega)$ is dominated by relatively narrow TA/LA peaks (hybridization at fixed $\omega$), while at higher frequencies the distribution broadens in $k$ space as disorder-induced attenuation becomes strong. [53, 35] In our continuum formalism, the corresponding broadening is captured in the imaginary part of the effective propagators $(-\omega^{2}+\tilde{C}_{\alpha}(\omega)k^{2})^{-1}$ appearing in Eq. (4).
S11. Rescaling procedure for the $R$-scan in Fig. 4
@@
Figure S2: (a) Real and imaginary parts of the effective shear modulus $G(\omega)$ from the HET-CPA analysis, plotted as $\mathrm{Re}[G(\omega)/G_{0}]$ (orange) and $\mathrm{Im}[G(\omega)/G_{0}]$ (blue), with $G_{0}=5.43~\mathrm{GPa}$. (b) Reduced vibrational density of states $g(\omega)/(A_{\mathrm{D}}\omega^{2})$ computed from $G(\omega)$ (magenta line) compared with INS data from Ref. [50] (filled circles).